OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (−1, 8)? y + 5 = 1(x + 2) y + 8 = −1(x − 1) y − 5 = −1(x − 2) y − 8 = 1(x − 2)

pooja195 (pooja195):

\[y-y1=m(x-x1)\] thats the point slop formula

pooja195 (pooja195):

ok first we need to find the slope so...\[\frac{ y2-y1 }{ x2-x1 }\] this is the slop formula \[\frac{ 8-5 }{ -1-2 }\]

pooja195 (pooja195):

can you solve that?

pooja195 (pooja195):

hello are you stuck?

pooja195 (pooja195):

\[\frac{ 3 }{ -3 }\] thats what you get when you solve the formula and when you simplfy you get -1 because -3/3=-1 now plug in the numbers to the slop formula i'll leave that part to you :3

pooja195 (pooja195):

alsp welcome to openstudy :)

pooja195 (pooja195):

*also

OpenStudy (anonymous):

the third one